OpenStudy (anonymous):

Please help! :) Two students in your class, Wilson and Alexis, are disputing a function. Wilson says that for the function, between x = -1 and x = 1, the average rate of change is 0. Alexis says that for the function, between x = -1 and x = 1, the graph goes up through a turning point, and then back down. Explain how Wilson and Alexis can both be correct, using complete sentences.

OpenStudy (anonymous):

Consider this graph... the rate of change between x=-1 and x=1 is 0 because they have the same y-value. the graph goes up to a turning point (at x=0) and then comes back down. 

OpenStudy (anonymous):

So the average rate of change is 0 because the y-value is the same?

OpenStudy (anonymous):

yes...rate of change is another name for a slope of a line. if you've studied lines and linear equations, you may remember that the slope is found to be \[m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] so if the y-values are the same for any two points, then the rate of change between those points will be 0. make sense?

OpenStudy (anonymous):

yes! so in other words Wilson and Alexis can both be correct because the y-value is the same?

OpenStudy (anonymous):

almost...that's really only half of it. Wilson can be correct if, for the function, f(-1) = f(1)...meaning the y-values are the same at x=-1 and x=1 Alexis can be corret if the graph, between x = -1 and x = 1 does what she claims. Do you know what kind of function has a graph like the one i drew? it would be useful to include that in your explanation...

OpenStudy (anonymous):

Quadratic?

OpenStudy (anonymous):

excellent! :) So if the function is quadratic opening down (a requirement for Alexis' claim) with the vertex occurring at x=0 (a requirement for Wilson's claim) then f(-1) = f(1) meaning the average rate of change is 0 between x=-1 and x=1. Since it opens down, it increases between x=-1 and x=0, turns around at x=0, and decreases from x=0 to x=1. :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you are quite welcome! :)